what is trigonometric expansions?

what is trigonometric expansions?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear nithin

Representation of power of trignometric ratio in terms trignometric ratio of multiple angel is known as trignometric expansions

 

for example.

cos6q = 1/32{cos 6q+6cos 4q+15cos 2q+10 }

or

cos7q = 1/64{cos 7q+7cos 5q+21cos 3q+35cos q  }

 

these expantion can bederived by puting cosx =(eix+e-ix)/2

  and expand it with the hepl of binomial .

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin